# More for Keys, Less for Values: Adaptive KV Cache Quantization

### v1

@@ 2.4 Total Energy of KV Cache (MSE Perspective)
Figure 3: Singular value distribution of key and value cache for the Llama 3.3 70B model on the C4 dataset. The x-axis represents the singular value indices, ordered from the 5th largest to the smallest, while the y-axis denotes the magnitude of the singular values. The shaded region indicates the minimum-to-maximum range across attention heads within each layer, and the solid lines represent the mean singular value magnitude at each index. This zoomed-in view reveals that, beyond the largest singular value (spectral norm), key matrices exhibit consistently higher singular values even in the lower ranges, underscoring their greater representational significance compared to value matrices. Consequently, key matrices should be quantized with less aggressive compression. See Figure 7 in Appendix C for the complete range.
Larger Frobenius Norms Require Higher Precision.
As we discussed in  2.1 empirical measurements across multiple transformer models reveal that key matrices have significantly larger Frobenius norms than value matrices. This disparity suggests that if key matrices are quantized too aggressively (e.g., with fewer bits), the MSE incurred could be disproportionately high. Formally, for two matrices $K$ (key) and $V$ (value), if
$\|K\|_{F}\;\gg\;\|V\|_{F},$
then allocating the same limited bit precision to both $K$ and $V$ would cause
$\|K-\widetilde{K}\|_{F}^{2}\;\gg\;\|V-\widetilde{V}\|_{F}^{2},$
since the error term for $K$ is amplified by its larger energy. Furthermore, errors introduced in earlier layers propagate through subsequent layers in a multi-layer transformer architecture, compounding the negative impact of coarse quantization.
3 Method
3.1 Adaptive KV Cache Quantization (KV-AdaQuant)
Building on the theoretical findings that matrices with higher spectral norms propagate quantization error more strongly and that larger Frobenius norms incur greater quantization error, along with empirical evidence indicating that key matrices exhibit higher norms than value matrices, a straightforward strategy emerges: quantize keys with higher precision and values with fewer bits. In particular, allocating 4-bit precision to keys and 2-bit precision to values effectively controls the overall mean squared error (Figure 4), reducing memory usage while mitigating error propagation through successive transformer layers.
(a) Mean4squared quantization error across layers for 2-bit quantization.
(b) Mean squared quantization error across layers for 4-bit quantization.
Figure 4: Quantization error (mean squared error) of key and value cache for the Llama 3.3 70B model on the C4 dataset across all transformer layers. The x-axis represents the layer index, and the y-axis indicates the mean squared quantization error. Key matrices exhibit consistently higher quantization error compared to value matrices, with larger discrepancies observed in lower-bit quantization.
3.2 Differences between KV-AdaQuant and Other Methods
The Hugging Face Transformers library implements a KV cache quantization technique that applies uniform precision to KV cache. Although straightforward, this approach can be suboptimal due to the distinct sensitivities of keys and values to quantization errors. Consequently, assigning the same bit-width to both key and value often introduces notable performance degradation. Furthermore, increasing the bit-width for value matrices offers minimal or no additional benefits, indicating that uniform quantization is not an ideal choice.
Other methods, such as SKVQ Duanmu et al. (2024) and QAQ Dong et al. (2024), also propose low-bit quantization for the KV cache, leveraging mixed-precision settings. SKVQ, for example, quantizes the key cache to 2 bits and the value cache to 1.5 bits, demonstrating minimal accuracy loss; however, it relies on fixed hyperparameter configurations that do not deeply analyze the key and value cache’ internal characteristics. QAQ similarly introduces quality-adaptive quantization to highlight the heightened sensitivity of key matrices to quantization errors, yet does not thoroughly investigate the underlying matrix properties, making generalization across diverse model architectures and tasks less certain.
By contrast, the proposed approach offers a more general framework of cache quantization that explicitly accounts for the distinct characteristics of key and value cache. This strategy reduces error propagation across transformer layers while achieving considerable memory savings. Additionally, it remains adaptable to further enhancements, such as calibration, grouping, and clustering, allowing the integration of complementary techniques to improve KV cache compression. This adaptability renders the framework orthogonal to existing methods, providing a more comprehensive solution to the challenges associated with KV cache quantization.
4 Experiments
A series of experiments is designed to investigate singular value patterns in KV cache and to demonstrate how these patterns can be leveraged to enhance practical quantization strategies. The experimental testbed consists of a diverse set of transformer models. Two distinct data setups are employed: one for analyzing the singular value distributions and another for evaluating the effectiveness of the proposed quantization approach.
4.1 Experiment Settings
Singular Value and Norm Analysis
This section examines structural differences between key and value cache by analyzing their singular values and norms. The process begins by feeding random batches (10 samples per batch) from the C4, MMLU, and GSM8K datasets into the model for each query, where every sample is padded to match the longest sequence to facilitate observations across diverse inputs. For each newly generated token (up to 1,000 tokens), the KV cache from each attention mechanism is collected, yielding matrices $\mathbf{K},\mathbf{V}\in\mathbb{R}^{d_{\text{head}}\times\text{seq\_len}}$ for each head and layer. Singular value decomposition is then performed offline in float32 precision using the PyTorch torch.linalg.svd operator, and the resulting singular values, along with the spectral norms ($\|\cdot\|_{2}$) and Frobenius norms ($\|\cdot\|_{F}$), are computed and averaged across all heads and layers to capture global patterns. This approach reveals how $\|\mathbf{K}\|_{2}$ and $\|\mathbf{V}\|_{2}$ evolve in deeper layers, sheds light on the total “energy” in these matrices, and clarifies the distribution of larger and smaller singular values, offering insights into their susceptibility to quantization errors.
Quantization Error
Hardware typically supports 2-bit, 4-bit, and 8-bit quantization, but in theory, any bit-width can balance compression and accuracy. We measure quantization error across bit-widths from 2 to 8 by quantizing and dequantizing key-value matrices, finding that MSE (and thus the Frobenius norm) behaves similarly across all bit-widths. As shown in Figure 5, the key matrix consistently exhibits higher error.
Bit-Width Configurations.
The effect of different bit-widths used to represent each entry in $\mathbf{K}$ and $\mathbf{V}$ is quantified. Uniform settings include bit-widths ranging from 2 to 8 bits for both keys and values (e.g., $(2,2)$, $(4,4)$, or $(8,8)$). Mixed-precision settings compare $(2,4)$ and $(4,2)$, assigning 2 bits to either values or keys and 4 bits to the other.
Figure 5: Mean squared error (MSE) of KV cache quantization for the Llama 3.3 70B model on the C4 dataset, using quantization bit-widths ranging from 2 to 8. The x-axis represents the quantization bit-width, while the y-axis shows the MSE on a logarithmic scale. A logarithmic scale is used to highlight differences at higher bit-widths, where MSE values decrease significantly and approach zero, particularly at 8-bit. Solid lines indicate the mean MSE across layers, while the shaded regions represent the min-max range of errors.
Quantization Procedure.
Hugging Face’s Optimum library is used for low-level CUDA kernels. No additional outlier or chunk-based grouping techniques are applied, so as to isolation of the effect of $\mathbf{K}$-vs-$\mathbf{V}$ bit allocation.
Error Metric: MSE.
To measure the quantization distortion, we record
$\text{MSE}(\mathbf{K},\widetilde{\mathbf{K}})\;=\;\frac{1}{d_{\text{head}}\,\times\,\text{seq\_len}}\,\|\mathbf{K}-\widetilde{\mathbf{K}}\|_{F}^{2},$
and likewise for $\mathbf{V}$. In addition to per layer and per head comparison, we average MSE across all heads, layers, and tokens in a batch. This approach yields a global view of how severely each bit-width compresses $\mathbf{K}$ vs. $\mathbf{V}$.
Experimental Protocol.
For each model and dataset subset (C4, MMLU, GSM8K):
  1. 1.
    Run a forward pass to produce unquantized KV caches in bfloat16.
  2. 2.
    Quantize to a chosen bit-width, then immediately dequantize to bfloat16.
  3. 3.
    Compare the dequantized version $\widetilde{\mathbf{K}}/\widetilde{\mathbf{V}}$ to the original $\mathbf{K}/\mathbf{V}$ in MSE terms.
  4. 4.
    Repeat for all bit-width settings of interest.
@@ 4.1 Experiment Settings
Table 1: Quantization error (mean ± std) for key and value cache at 2-bit, 3-bit, and 4-bit quantization ($K_{i}$ and $V_{i}$) across different datasets.
@@
Downstream Evaluation.
While MSE is a reasonable proxy for overall distortion, the ultimate question is whether the quantized KV cache preserves model quality on realistic tasks. We therefore evaluate the final model performance on GSM8K in two prompting modes: 1-shot and 8-shot prompting using 4 quantization settings:
  • $\mathbf{K}_{\mathbf{(2bit)}}\mathbf{V}_{\mathbf{(2bit)}}$: Aggressive uniform quantization.
  • $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(4bit)}}$: More conservative uniform quantization.
  • $\mathbf{K}_{\mathbf{(2bit)}}\mathbf{V}_{\mathbf{(4bit)}}$ vs. $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(2bit)}}$: Mixed-precision assignments, giving 2 bits to either $\mathbf{K}$ or $\mathbf{V}$ and 4 bits to the other.
4.2 Experiment Results
Table 2: GSM8K 1-shot and 8-shot performance of various models under different key-value quantization settings. $\mathbf{K}_{i}\mathbf{V}_{j}$ denotes $i$-bit quantization for the K cache and $j$-bit for the V cache. The $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(2bit)}}$ column, highlighted below, represents our main finding.

### v2

@@ 3.2 Key-Driven Quantization
Figure 3: Singular value distribution of key and value activations for Llama 3.3-70B on C4. The x-axis represents the singular value indices, ordered from the 5th largest to the smallest, while the y-axis denotes the magnitude of the singular values. The shaded region indicates the minimum-to-maximum range across attention heads within each layer, and the solid lines represent the mean singular value magnitude at each index. This zoomed-in view reveals that, beyond the largest singular value (spectral norm), key matrices exhibit consistently higher singular values even in the lower ranges, underscoring their greater representational significance compared to value matrices. For the complete range, see Figure 5 in Appendix E.
Appendix 3 and  4 establishes a norm-dependent upper bound on the quantization error of an arbitrary matrix $M$, and Appendix C.6 applies this bound to show that key caches incur proportionally larger errors than value caches.
To empirically validate the theoretical predictions, we examined the singular value spectra of the K anc V caches across a broad set of models, up to 70B, by extracting every attention head in every layer (Section 4.1). Figure 3 displays the singular values from the fifth largest onward, revealing a systematic separation between keys and values; the complete spectra in Appendix E (Figure 5) show the same trend across all modes. Aggregate norms reinforce this observation: Figures 7 and 7 plot the layer-wise Frobenius and spectral norms of the $K\!V$ cache in Llama-3.3-70B, consistently indicating larger magnitudes for keys than for values.
4 Experimental Results
4.1 Experiment Setup
We conduct two experiments: (i) a quantization-error analysis and (ii) a downstream evaluation. For the quantization-error analysis, we use PyTorch’s weight-packing quantization to measure reconstruction error. For the downstream evaluation, we employ hardware-aligned quantization via Hugging Face’s Optimum Quanto library as back-end (token-wise, 2/4-bit quantization) on Llama-3.2-1B, Llama-3.1-8B [38], Phi-4-14B [39] and DeepSeek-R1-Qwen-14B, and the Half-Quadratic Quantization (HQQ) [40] back-end (channel-wise, 1–8-bit quantization) on Llama-3.1-8B, Llama-3.2-1B, Llama-3.2-3B, and Qwen3-0.6B/1.7B/4B/8B [41]. The quantization-error sweep additionally includes Phi-3-Medium-128K [42], Llama-3.1-Nemotron-70B [43], Mistral-0.3-7B [8], Qwen-2.5-14B and Llama-3.3-70B. See Appendix G for full details on compute resources, software tools, and reproducibility.
@@ 4.1 Experiment Setup
Table 1: Quantization errors for KV cache under various configurations. The quantization error is expressed as mean ± std. We use 2-bit, 3-bit, and 4-bit quantization for both $K_{i}$ and $V_{i}$ across three datasets.
@@
4.2 Quantization Error
In several studies (Table 3 of KIVI[29], Table 8 of GEAR[44], and Table 4 of RotateKV[45]), quantized KV cache representations have been shown to outperform their BF16 full-precision counterparts, an outcome that may seem counterintuitive. Taking BF16 as the gold standard, we attribute this phenomenon to the implicit regularization effect introduced by quantization: by discretizing activation values, quantization injects a small, structured noise into the model’s intermediate representations. Much like dropout or weight decay, this noise can smooth overly sharp decision boundaries and mitigate over-fitting, thereby improving generalization on held-out tasks. To isolate and measure this effect, our first experiment focuses on characterizing the quantization error in the KV cache, without grouping tokens, or using residual tokens in each quantization turn.
In addition to per layer and per head comparison, we average MSE across all heads, layers, and tokens in a batch. This approach yields a global view of how severely each bit-width compresses $\mathbf{K}$ vs. $\mathbf{V}$.
The evaluation uses ten random sequences from C4 [46], MMLU [47], and GSM8K [48], padded to the longest length and generated autoregressively for up to 1 000 tokens. Table 1 presents 2-, 3-, and 4-bit errors for seven representative models (Llama3, Phi4 [39], Mistral, Qwen, DeepSeek-Distilled, and others); the complete results appear in Appendix F. Figure 8 plots the full error curve for Llama-3.3-70B on C4, while Figure 10 shows the per-layer breakdown for Llama-3.1-8B.
4.3 Evaluation
Experiments in Section 4.2 established that K cache incur larger MSE than V cache at every bit-width, suggesting that $\mathbf{K}$ contains proportionally richer information. However, ultimately model quality is measured by downstream metrics such as accuracy rather than reconstruction error. Our second experiments therefore measures how the observed error asymmetry translates into task performance under realistic, deployment-oriented constraints.
KV cache quantization methods rarely de-quantize the entire cache after every generated token because the associated memory traffic defeats the purpose of quantization. Instead, they maintain a residual buffer that holds the most recent tokens in full-precision and flushes it en bloc once full. We replicate this strategy by (i) reserving a fixed-size residual cache of 64 tokens and (ii) grouping activations into 128-element blocks—settings adopted by popular open-source implementations such as KIVI, Flash-Decoding [49] and Marlin [50]. Two hardware-aligned quantization kernels were selected to span the design space:
  • Quanto: token-wise (per-row) quantization applied to the key and value caches, supporting $\{2,4\}$-bit precision.
  • HQQ: channel-wise (per-column) quantization applied to the key and value caches, supporting $\{1,2,4,6,8\}$-bit precision.
For downstream evaluation, we employ three generative datasets that require the model to produce extended outputs rather than select among fixed choices: the GSM8K, the CoQA [51] conversational question answering dataset, and the EQ-Bench [52] long-form structured QA suite.
Results.
Tables 2 and 3 corroborate the norm-based hypothesis by showing that downstream accuracy is far more sensitive to the precision assigned to the key cache than to the value cache.
  • Quanto backend. Across four models and two prompt lengths, allocating $4$ bits to keys and $2$ bits to values ($\mathbf{K}_{4}\mathbf{V}_{2}$) recovers on average $94\%$ of the full-precision score, whereas the inverted scheme ($\mathbf{K}_{2}\mathbf{V}_{4}$) loses up to $30$ percentage points (pp). The margin widens with model capacity: in 1-shot GSM8K, $\mathbf{K}_{4}\mathbf{V}_{2}$ outperforms $\mathbf{K}_{2}\mathbf{V}_{4}$ by $+0.30$ pp on Llama-3.2-1B and by $+0.16$ pp on Phi-4-14B. Notably, $\mathbf{K}_{4}\mathbf{V}_{2}$ attains almost the same accuracy as the symmetric $4$-bit baseline.
  • HQQ backend. The broader sweep over $\{1,2,4,6,8\}$ bits and three datasets paints the same picture. On GSM8K, granting even a single extra bit to keys yields double-digit gains; e.g., Qwen3-8B improves from $0.08$ to $0.54$ accuracy ($+46$ pp) when the extra bit is moved from values to keys. Gaps persist on EQ-Bench and CoQA, where key-centric quantization delivers up to $+62$ pp (Qwen3-0.6B, EQ-Parseable, $1$-bit) and never underperforms the converse setting. Advantages taper but do not vanish at higher precisions: at $6$ bits the key-first policy still adds $7\!-\!12$ pp for 8–14B models.
Table 2: KV cache quantization on GSM8K 1-shot and 8-shot performance employing Optimum Quanto backend. $\mathbf{K}_{i}\mathbf{V}_{j}$ denotes $i$-bit quantization for the K cache and $j$-bit for the V cache using Quanto backend. The $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(2bit)}}$ column, highlighted below, shows the optimal bit allocation setting.